OpenStudy (shainabeans):

find cos x/2, given cosx=1/4, with 0

OpenStudy (callisto):

cosx = 2cos^2 (x/2) -1 1/4 = 2cos^2 (x/2) -1 5/4 = 2cos^2 (x/2) 5/2 = cos^2 (x/2) \[cos (x/2) = \sqrt{5/2} =? \]

OpenStudy (anonymous):

true ;

OpenStudy (callisto):

Actually I've given the answer :| Silly me to type that question mark!!!!

OpenStudy (anonymous):

lool it causes me no problems marks